# Probing modified Hawking evaporation with gravitational waves from the primordial black hole dominated universe

### v1

@@ IV Discussion and conclusions
We considered two models for the memory burden inspired by Ref. [19]. One where the modification only involves a constant entropy factor (case mb1) and another where the entropy is a function of the evaporating PBH mass (case mb2). The relevant difference between these two cases is that in the mb2 case, the final burst of the PBHs happens faster than in the mb1 and standard cases. The two models are very similar otherwise, leading to a much longer stage of PBH domination. However, the faster final evaporation changes the slope of the GW spectrum of GWs induced by PBH number density fluctuations, giving a unique signature of the model mb2.
@@
It is worth noting that our results summarized above apply to more general modifications of Hawking evaporation. In particular, since the modifications of Hawking evaporation are a local effect (i.e., related to the individual black holes), it will not affect the $\beta$ dependence we obtain in our calculations. For instance, the peak frequency $f_{\rm uv}$ (24) only depends on $M_{\rm f}$ and the peak amplitude (34) will be proportional to $\beta^{16/3}$. We expect that $f_{\rm uv}$ will be bound to frequencies $\gtrsim\rm Hz$ due to BBN bounds on the evaporation temperature. Moreover, if the modification affects the time dependence of the evaporating mass (i.e., not only extending it by a constant factor), the slope of the GW spectrum will be sensitive to such modification. Lastly, it is interesting to note that other effects, such as PBH spin and additional degrees of freedom (that is, increasing the value of $g_{H}$), only accelerate evaporation slightly. For instance, the evaporation time of highly spinning PBHs is about a factor $1/2$ smaller than that in the zero spin case [63, 64]. The impact on the induced GW spectrum is thus not significant [28].
@@
As extensions of this work, it would be interesting to study the impact of the memory burden on the PBH remnants as dark matter [65] (see Refs. [66, 67, 68, 69, 70] for details on black hole remnants) and their unique GW signature [71], which should require smaller PBH masses for larger values of $n$. It would also be interesting to study how the spectrum of induced GWs from primordial adiabatic fluctuations, such as the ones measured by CMB observations, is enhanced due to the PBH-dominated phase as in Ref. [25]. Our expectation is that it should be similar to the PBH number density fluctuations studied in this paper, although it might depend on whether one employs a cut-off for scales that become non-linear at evaporation [25]. It would also be interesting to explore the effect of an extended mass function in more detail [25, 30]. Lastly, it would be interesting to explore the modifications to the GWs from Hawking evaporation [72, 63, 37, 38, 73, 60, 74, 75]. It should also be noted that one known issue of the GWs induced by PBH density fluctuations from Refs. [26, 27], is that some scales enter the non-linear regime before evaporation. However, curvature perturbations (which is the main source of induced GWs) remain within the regime of validity of perturbation theory, and so the calculations of the induced GW spectrum can be understood as an (optimistic) estimate. We checked in App. B that including the memory burden effect worsens this issue. Nevertheless, the main point of this paper is to estimate the impact of the memory burden effect on the PBH reheating scenario, which should not be affected by non-linear effects. To derive more accurate estimates, one requires more sophisticated numerical simulations.
@@ Acknowledgements.
We would like to thank I. Masina and V. Takhistov for useful feedback. SB is supported by the STFC under grant ST/X000753/1. GD, AG, and JT are supported by the DFG under the Emmy-Noether program, project number 496592360.
@@ Appendix B Evolution of curvature fluctuations induced by PBH number density fluctuations
In this case, the suppression is $k$-dependent because we have that $\Gamma\to\infty$ as $t\to t_{\rm eva}$ and eventually all $k$’s decouple.
@@
$\frac{\delta\rho_{\rm PBH}}{\rho_{\rm PBH}}(k_{\rm uv})\Bigg{|}_{\rm eva}\approx\sqrt{\frac{2}{3}}\left(\frac{3}{4}\right)^{\frac{1}{6\alpha}}\,q\,\gamma^{2/3}\beta^{4/3}\left(\frac{\alpha}{\sqrt{3\alpha-1}}\right)^{-\frac{1}{3\alpha}}\frac{e^{\frac{(6\alpha-1)(17+16n)}{6\alpha}}}{\left({\alpha}/{q}\right)^{\frac{2}{3}-\frac{1}{9\alpha}}}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{\frac{2(6\alpha-1)(1+n)}{9\alpha}}\,.$ (49)
@@
$\frac{\delta\rho_{\rm PBH}}{\rho_{\rm PBH}}(k_{\rm uv})\bigg{|}^{\rm max}_{\rm eva}\approx 25\left(\frac{3}{4}\right)^{\frac{1}{6\alpha}}\frac{e^{2n}}{q^{11/12}\alpha^{1/12}}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{\frac{1+n}{6}}\,,$ (50)
@@
$\frac{\delta\rho_{\rm PBH}}{\rho_{\rm PBH}}(k_{\rm uv},n=1)\bigg{|}^{\rm max}_{\rm eva}\approx 340\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{\frac{1}{3}}\,.$ (51)
@@ Appendix C Induced Gravitational Wave Integrals
$\Omega_{\rm GW}=\frac{k^{2}}{12\mathcal{H}^{2}}\overline{\mathcal{P}_{h}}$ $(k,\tau)$ (52)
@@
×$\times\int_{0}^{\infty}\frac{d\tilde{x}}{\tilde{x}+x_{\rm eva}/2}\sin(x_{1}-\tilde{x})\sin\left(c_{s}u\tilde{x}\right)\sin\left(c_{s}v\tilde{x}\right)$ (54)
@@
Figure 7: We compare the numerical GW spectral densities $\Omega_{\rm GW}(f)$ (solid lines), normalized by $\beta^{16/3}$, with parameters $M_{\rm PBH,f}=10^{4}$\mathrm{g}$$ and $n=\frac{1}{2}$ for both scenarios, mb1 and mb2, with the respective analytical result for the resonant contribution Eq. (29) (dashed lines). As can be seen, the analytical approximation captures the peak very well, with deviations due to the IR tail setting in at around $f/f_{\rm uv}\sim 10^{-3}$ (for this case of $n=\frac{1}{2}$), and above the resonant scale $f_{\rm res}=2c_{s}f_{\rm uv}$ in the UV. We found that the amplitude of the numerical solution is larger than the analytical one by a factor of 2 (likely stemming from the various approximations). We corrected the factor in the plot to show the good agreement of the spectral shape.

### v2

@@ IV Discussion and conclusions
We considered two models for the memory burden inspired by Ref. [22]. One where the modification only involves a constant entropy factor (case mb1) and another where the entropy is a function of the evaporating PBH mass (case mb2). The relevant difference between these two cases is that in the mb2 case, the final burst of the PBHs happens faster than in the mb1 and the standard case. The two models are very similar otherwise, leading to a much longer stage of PBH domination. However, the faster final evaporation changes the slope of the GW spectrum of GWs induced by PBH number density fluctuations, giving a unique signature of the model mb2.
@@
It is worth noting that our results summarized above apply to more general modifications of Hawking evaporation. In particular, since the modifications of Hawking evaporation are a local effect (i.e., related to the individual black holes), it will not affect the $\beta$ dependence we obtain in our calculations. For instance, the peak frequency $f_{\rm uv}$ (24) only depends on $M_{\rm f}$ and the peak amplitude (34) will be proportional to $\beta^{16/3}$. We expect that $f_{\rm uv}$ will be bound to frequencies $\gtrsim\rm Hz$ due to BBN bounds on the evaporation temperature. Moreover, if the modification affects the time dependence of the evaporating mass (i.e., not only extending it by a constant factor), the slope of the GW spectrum will be sensitive to such modification. Lastly, it is interesting to note that other effects, such as PBH spin and additional degrees of freedom (that is, increasing the value of $g_{H}$), only accelerate evaporation slightly. For instance, the evaporation time of highly spinning PBHs is about a factor $1/2$ smaller than that in the zero spin case [80, 81]. The impact on the induced GW spectrum is thus not significant [33, 38].
@@
As extensions of this work, it would be interesting to study the impact of the memory burden on the PBH remnants as dark matter [82] (see Refs. [83, 84, 85, 86, 87] for details on black hole remnants) and their unique GW signature [88], which should require smaller PBH masses for larger values of $n$. It would also be interesting to study how the spectrum of induced GWs from primordial adiabatic fluctuations, such as the ones measured by CMB observations, is enhanced due to the PBH-dominated phase as in Ref. [30] (see also Refs.[37, 38]). Our expectation is that it should be similar to the PBH number density fluctuations studied in this paper, although it might depend on whether one employs a cut-off for scales that become non-linear at evaporation [30]. It would also be interesting to explore the effect of an extended mass function in more detail [30, 35] and study the effect of non-Gaussianities and PBH clustering in conjuction with the memory burden effect [89, 90]. Another interesting aspect to consider is the fact that black hole thermodynamics depends on the underlying theory of gravity itself. In particular, the definition of the black hole entropy (2) changes in theories of modified and quantum gravity [91, 92, 93]. The induced GWs from the PBH reheating scenario could provide a unique opportunity to probe such effects, which are otherwise hard to access. Lastly, it would be interesting to explore the modifications to the GWs from Hawking evaporation [94, 80, 44, 45, 95, 77, 96, 97]. It should also be noted that one known issue of the GWs induced by PBH density fluctuations from Refs. [31, 32], is that some scales enter the non-linear regime before evaporation. However, curvature perturbations (which is the main source of induced GWs) remain within the regime of validity of perturbation theory, and so the calculations of the induced GW spectrum can be understood as an (optimistic) estimate. We checked in App. B that including the memory burden effect worsens this issue. Nevertheless, the main point of this paper is to estimate the impact of the memory burden effect on the PBH reheating scenario, which should not be affected by non-linear effects. To derive more accurate estimates, one requires more sophisticated numerical simulations.
@@ Acknowledgements.
We would like to thank I. Masina and V. Takhistov for useful feedback. We would also like to thank R. Haque for pointing out a typo in a previous version of the manuscript. SB is supported by the STFC under grant ST/X000753/1. GD, AG, and JT are supported by the DFG under the Emmy-Noether program, project number 496592360. GD is also supported by the JSPS KAKENHI grant No. JP24K00624.
@@ Appendix B Evolution of curvature fluctuations induced by PBH number density fluctuations
In this case, the suppression is $k$-dependent because we have that $\Gamma\to\infty$ as $t\to t_{\rm eva}$ and eventually all $k$’s decouple. Let us note here that, in the case of a monochromatic PBH mass function, the instantaneous transition approximation with the suppression factor (46) has been shown to be in good agreement with numerical results [30], especially for the very small scales we are most interested in. In the case of a broader mass function the transition would be more gradual, and the amplitude of the induced GWs would be slightly suppressed. The effects of an extended mass function and a gradual transition have been discussed in [35, 64, 36], respectively.
@@
$\frac{\delta\rho_{\rm PBH}}{\rho_{\rm PBH}}(k_{\rm uv})\Bigg{|}_{\rm eva}\approx\sqrt{\frac{2}{3}}\left(\frac{3}{4}\right)^{\frac{1}{6\alpha}}\,q\,\gamma^{2/3}\beta^{4/3}\left(\frac{\alpha}{\sqrt{3\alpha-1}}\right)^{-\frac{1}{3\alpha}}\frac{e^{\frac{(6\alpha-1)(17+16n)}{6\alpha}}}{\alpha^{\frac{2}{3}-\frac{1}{9\alpha}}}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{\frac{2(6\alpha-1)(1+n)}{9\alpha}}\,.$ (49)
@@
$\frac{\delta\rho_{\rm PBH}}{\rho_{\rm PBH}}(k_{\rm uv})\bigg{|}^{\rm max}_{\rm eva}\approx 25\left(\frac{3}{4}\right)^{\frac{1}{6\alpha}}\frac{e^{2n}}{\alpha^{1/12}}\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{\frac{1+n}{6}}\,,$ (50)
@@
$\frac{\delta\rho_{\rm PBH}}{\rho_{\rm PBH}}(k_{\rm uv},n=1)\bigg{|}^{\rm max}_{\rm eva}\approx 160\,\left(\frac{M_{\rm mb}}{1\,\rm g}\right)^{\frac{1}{3}}\,.$ (51)
@@ Appendix C Induced Gravitational Wave Integrals
$\Omega_{\rm GW}=\frac{k^{2}}{12\mathcal{H}^{2}}\overline{\mathcal{P}_{h}}$ $(k,\tau)\,,$ (52)
@@
×$\times\int_{0}^{\infty}\frac{d\tilde{x}}{\tilde{x}+x_{\rm eva}/2}\sin(x_{1}-\tilde{x})\sin\left(c_{s}u\tilde{x}\right)\sin\left(c_{s}v\tilde{x}\right)\,,$ (54)
@@
Figure 7: We compare the numerical GW spectral densities $\Omega_{\rm GW}(f)$ (solid lines), normalized by $\beta^{16/3}$, with parameters $M_{\rm PBH,f}=10^{4}$\mathrm{g}$$ and $n=\frac{1}{2}$ for both scenarios, mb1 and mb2, with the respective analytical result for the resonant contribution Eq. (29) (dashed lines). As can be seen, the analytical approximation captures the peak very well, with deviations due to the IR tail setting in at around $f/f_{\rm uv}\sim 10^{-3}$ (for this case of $n=\frac{1}{2}$), and above the resonant scale $f_{\rm res}=2c_{s}f_{\rm uv}$ in the UV. We found that the amplitude of the numerical solution is larger than the analytical approximation (29) for the resonant peak by a factor of 2, stemming from the contribution of the non-divergent terms resulting from (54). We corrected the factor in the plot to show the good agreement of the spectral shape.